# Generalized Jeffreys’s approximate objective

### v2

@@ Abstract
Concerns about the misuse and misinterpretation of $p$-values and statistical significance have motivated alternatives for quantifying evidence. We define a generalized form of Jeffreys’s approximate objective Bayes factor ($\mathit{eJAB}$), a one-line calculation that is a function of the $p$-value, sample size, and parameter dimension. We establish conditions under which $\mathit{eJAB}$ is model-selection consistent and verify them for ten common statistical tests. We assess finite-sample accuracy by comparing $\mathit{eJAB}$ with Markov chain Monte Carlo computed Bayes factors in 12 simulation studies. We then apply $\mathit{eJAB}$ to 71,126 results from ClinicalTrials.gov (CTG) and find that the proportion of findings with $p\text{-value}\leqslant\alpha$ yet $\mathit{eJAB}_{01}>1$ closely tracks the significance level $\alpha$, suggesting that such contradictions are pointing to the type I errors. We catalog 4,088 such candidate type I errors and provide details for 131 with reported $p$-value $\leqslant.01$. We also identify 487 instances of the Jeffreys–Lindley paradox. Finally, we estimate that 75% (6%) of clinical trial plans from CTG set $\alpha\geqslant.05$ as the target evidence threshold, and that 35.5% (0.22%) of results significant at $\alpha=.05$ correspond to evidence that is no stronger than anecdotal under $\mathit{eJAB}$.
@@
Concerns about the replicability of scientific research—stemming in part from the overreliance on null-hypothesis significance testing (NHST) at traditional $p$-value thresholds [1]—have prompted calls for more stringent evidence criteria [2, 3], more judicious interpretation [4], and alternative measures of evidence [5, 6, 7, 8]. The Bayes factor—the ratio of marginal likelihoods under competing hypotheses—can provide evidence for either the null or the alternative [9, 10, 11, 12, 13, 14]. It represents how data update the prior model odds to form the posterior model odds. The evidential behavior of both $p$-values and Bayes factors has been analyzed by stochastic ordering in sample size and effect size, with implications for the Jeffreys–Lindley paradox [15, 16]. Relatedly, emphasizing effect sizes rather than $p$-value significance thresholds, a quadratic exponential relationship between the Bayes factor and the separation of credible intervals is described in [17]. However, computation remains a practical hurdle: Bayes factors require numerical integration over prior distributions, often via Markov chain Monte Carlo (MCMC) with model-specific tuning [18, 19, 20]. Although these computational challenges are not insurmountable, a simple Bayes factor calculation that uses only the $p$-value, sample size, and parameter dimension has greater applicability to situations where the full dataset is unavailable.
@@
Several integral-free methods have been proposed to reduce computational complexity. For example, certain Bayes factors, with informed priors on a focal parameter, yield fast approximations and are asymptotically equivalent to Jeffreys’s approximate objective Bayes factor (JAB) [21, 22, 23, 24]. The Bayesian information criterion (BIC) is related to Bayes factors for model comparison [14, 25]. Minimum-Bayes-factor calibrations provide conservative evidence assessments when only a $p$-value is available [26]. Bayes factors based on parametric test statistics are developed by [27], with model-selection consistency under the prior predictive and connections to BIC examined in [28]; extensions to nonparametric statistics appear in [29]. A proposed $3p\sqrt{n}$ rule, based on a piecewise approximation to the $\chi^{2}_{1}$ quantile function, facilitates JAB in the scalar-parameter case [30]. We define $\mathit{eJAB}$, a generalization of JAB that incorporates a finite-sample correction and explicitly accounts for parameter dimension; it is a one-line calculation that is model-selection consistent under broad conditions. eJAB reduces to JAB when the parameter under test is a scalar and as $n\rightarrow\infty$.
@@
Our work differs from [25] in its consideration of $p$-values beyond those from the likelihood ratio test and in our explicit statement of conditions for model-selection consistency. For the latter, [25] demonstrates asymptotic equivalence to the BIC so that model-selection consistency arises under the conditions specified in [31]. It differs from [23, 24, 30] primarily in our development of a theoretical justification.
@@
Our contributions are thus threefold. First, we define an extension of JAB, $\mathit{eJAB}$, which generalizes to multidimensional parameters by incorporating the parameter dimension. Our approach differs from BIC-based approximations by accommodating $p$-values beyond those from likelihood-ratio tests, and from other methods by establishing a model-selection consistency justification under regularity conditions. Second, we assess the finite-sample performance of $\mathit{eJAB}$ across ten common parametric and nonparametric tests. The results suggest that $\mathit{eJAB}$ is a broadly applicable and reliable approach to approximate the Bayes factor. Third, we apply $\mathit{eJAB}$ to 71,126 reported findings from https://ClinicalTrials.gov (CTG), accessed on August 7th, 2025, flagging 4,088 candidate type I errors. We find that the distribution of statistical evidence is uniform across five clinical trial phases for both primary and secondary outcomes in CTG and identify 487 instances of the Jeffreys–Lindley paradox.
@@ 1 Methodology
where $q$ is the dimension of the parameter vector $\boldsymbol{\theta}$, $n$ is the sample size, $Q_{\chi^{2}_{q}}(\cdot)$ is the quantile function of the chi-squared distribution with $q$ degrees of freedom (df), and $p$ is the $p$-value from an NHST. The coherence condition specifies that $\mathit{eJAB}_{10}=1\ \!/\ \!\mathit{eJAB}_{01}$.
@@
Assuming the nuisance-parameter priors for the nested models satisfy the usual conditioning constraint—i.e., the prior under $\mathcal{H}_{0}$ equals the conditional prior under $\mathcal{H}_{1}$ given the null constraints—the Bayes factor can be expressed as the Savage–Dickey density ratio [32],
@@ 2 Simulation Studies
We used $p$-values from the $t$-test for items 1 and 2; the $F$-test for items 4 and 5; the Wald test for items 3 and 7; and the asymptotic null distribution for item 6. The $p$-values for 8-10 are based on asymptotic distributions. The effective sample size $n$ was taken as the number of uncensored observations for survival analysis and as the number of independent observations (i.e., the product of the number of participants and one less than the number of conditions) for rANOVA [33, 34]. For the other designs, $n$ always refers to the total number of observations or counts. In particular, we tested the harmonic mean of group sizes for item 1, as in JAB [30], but do not demonstrate it here. Among the tests with multiple df ($q>1$) are items 4 through 6, and 10.
Figure 1: Sampling distributions of $D_{n}=-\sqrt{n}\cdot p\ln{p}$ under $\mathcal{H}_{1}$ for ten statistical tests. Panel labels: ttest ($t$-test), lm (linear regression), glm (generalized linear regression), anova (analysis of variance), ranova (repeated-measures ANOVA), chisq (chi-squared test), coxph (Cox proportional-hazards regression), wilcox (Wilcoxon signed-rank test), Utest (Mann–Whitney test), and Htest (Kruskal–Wallis test).
Figure 2: Comparison of the natural logarithmic $\mathit{eJAB}_{10}$ with alternative Bayes-factor computations at a fixed sample size. The $y$-axis shows the MCMC-based Bayes factor for all panels, except for the two chi-squared tests, which use the test-statistic Bayes factor instead. The red text indicates the effective sample size $n$, except for the Cox model, which shows the number of observations. The dashed lines indicate $\mathit{eJAB}_{10}$ of 1/3 and 3, representing moderate evidence against and for an effect, respectively.
@@ 2.1 Verifying the Regularity Conditions
Next, we computed $D_{n}$ of R2 using the $p$-values and sample sizes from the simulated data, generated with medium effect sizes, across the ten designs, plus conditional logistic regression used in our application. In this subsection, items 4 and 5 ($\rho=.9$) each have 41=34-1=3 df; item 6 has $(3-1)\times(3-1)=4$ df under a multinomial design; item 7 represents a semi-parametric Cox model; item 10 has 51=45-1=4 df. Due to randomness, the $x$-axis has to display the total number of observations or pairs, rather than the number of uncensored observations for item 7, or the number of matched pairs for conditional logistic regression, as required in $D_{n}$ and $\mathit{eJAB}$. Condition R2 appears to hold for all fixed alternatives we examined (Fig. 1). For local alternatives of the form $\theta^{*}=\theta_{0}+c/n^{\kappa}$, R2 does not hold at $\kappa=1/2$ but does hold at $\kappa=1/4$ (Appendix B, Fig. 5A).
@@ 2.2 Accuracy of the Bayes-Factor Approximation
Pairs of $(\ln\mathit{eJAB}_{10},\ \ln\mathit{BF}_{10})$ are color- and shape-coded by the $p$-value. Overall, $\mathit{eJAB}$ tracks the MCMC-based BFs closely (Fig. 2). Some inaccuracy, $\ln\mathit{eJAB}_{10}<\ln\mathit{BF}_{10}$, appears for (binary) logistic regression when evidence favors $\mathcal{H}_{1}$. For chi-squared tests, agreement with test-statistic BFs is excellent when evidence favors $\mathcal{H}_{1}$. As evidence shifts toward $\mathcal{H}_{0}$, $\mathit{eJAB}_{10}$ decreases toward zero, as expected under model-selection consistency, whereas the test-statistic BF is inherently bounded below by 1. Additional comparisons with Dirichlet BFs for chi-squared tests are provided in Appendix B. The two Bayes factors reach broad agreement, although differences persist at larger $n$ (Appendix B, Fig. 6). At $n=1,500$, under $\mathcal{H}_{0}$, the Dirichlet $\mathit{BF}_{10}<1/3$ is almost always true, whereas $\mathit{eJAB}_{10}<3$ is mostly observed; under $\mathcal{H}_{1}$, $\mathit{eJAB}_{10}$ more readily identifies an effect. Replacing $\mathit{eJAB}$ with the BIC approximation to the BF performs well when $q=1$, but marked severe discrepancies for $q>1$, indicating that $\mathit{eJAB}$ better approximates BF for ANOVA, rANOVA, and chi-squared tests (Appendix B, Fig. 7).
@@ 2.3 Bayesian Hypothesis Tests Using Nonparametric Statistics
3 Reevaluating Statistical Evidence in 71,126 Clinical Trial Findings with eJAB
@@
Our objective is to characterize the overall strength of evidence in 71,126 CTG findings and to identify potential type I errors. These results span 43 MeSH categories (e.g., neoplasms, cardiovascular diseases, and autoimmune diseases) and encompass nine statistical tests: one-sample $t$-test, two-sample $t$-test, linear regression, logistic regression, ANOVA, conditional logistic regression, Wilcoxon signed-rank test, Mann–Whitney test, and Kruskal–Wallis test. We verify that regularity condition R2 holds for conditional logistic regression (Appendix B, Fig. 5B).
@@
We therefore classify findings with $p\leqslant\alpha$ and $\mathit{eJAB}_{01}>1$ as candidate type I errors at the significance level $\alpha$. Among 30,790 results significant at $\alpha=.05$, we catalog 4,088 candidates. Although lower thresholds have been recommended in the literature [2, 3], $\alpha=.05$ remains germane in CTG: in a random sample of 52 trial protocols, 75% (SE 6%) explicitly set $\alpha\geqslant.05$ as the target evidential threshold. We list all 4,088 candidates on the Open Science Framework at https://osf.io/qeprj/ (see ‘Supplementary Information’) and provide an annotated subset of 131 cases with $p\leqslant.01$ and $\mathit{eJAB}_{01}>1$, cross-referenced to CTG and PubMed.
@@
We also examine the proportion of all $p\leqslant.05$ and $\mathit{eJAB}_{01}\geqslant K$ for $K\in(0,300)$; at $K=1$, this proportion recovers the candidate type I error at $\alpha=.05$. The inset highlights Jeffreys–Lindley paradox cases (Fig. 3C). We identify 487 such instances in the CTG data. Additional results stratified by MeSH category appear in Appendix B (Figs. 8-10).
@@
Figure 3: Evidence in 71,126 results. (A-B) Proportion of analyses with $p\leqslant\alpha$ and $\mathit{eJAB}_{01}>1$ as candidates type I errors, with $p=.05$ excluded in (B). The red dashed line marks the overall share with $\mathit{eJAB}_{01}>1$ as $\alpha\rightarrow 1$. (C) Proportion of $p\leqslant.05$ with Bayes-factor threshold $K$. The inset highlights the Jeffreys–Lindley paradox region. (D) Scatterplot of $\ln\mathit{eJAB}_{01}$ versus $p$-value, with an inset zoom for $p\leqslant.05$. Shaded bands denote evidence categories. (E) Posterior distribution of the proportion of $p\leqslant.05$ and $\mathit{eJAB}_{01}>1/3$. The red dashed line indicates the posterior mean. (F) Density of sample sizes for all $p\leqslant.05$, stratified by Bayes-factor category.
@@
We examine the distributions of $\ln\mathit{eJAB}_{01}$ and $\ln{p}$ by trial phase for primary and secondary outcomes (Fig. 4A-B). For primary outcomes, there is no discernible trend toward stronger evidence from early phase to phase 4. The distributions of $\ln{p}$ shift downward at phase 3, consistent with the markedly larger average sample sizes at that phase. For secondary outcomes, we again observe no phase-wise increase in evidence, though variability in both measures is higher in phase 3. Plotting $p$ and sample size by phase shows that changes in $p$ track changes in sample size (Fig. 12 in Appendix B).
To better visualize trends in evidence across phases, we fit a linear mixed model with a fixed intercept and trial-specific random effects to each of $\ln\mathit{eJAB}_{01}$ and $\ln{p}$, and then examine residuals re-centered by the estimated intercept (Fig. 11). The adjusted measures are essentially uniform across phases (Fig. 4C-D). For primary outcomes, most adjusted $p$-values lie near .05 and above .01. Using $\mathit{eJAB}$, this interpretation corresponds to ‘anecdotal evidence’ or ‘evidence that is barely worth mentioning’ for both primary and secondary outcomes [42, p. 105]. Comparing primary with secondary outcomes, we note a slight and phase-invariant shift toward greater evidence for the alternative in primary outcomes.
Figure 4: Measures of evidence by clinical trial phase. (A-B) Distributions of $\ln\mathit{eJAB}_{01}$ and $\ln{p}$ by phase for primary (A) and secondary (B) outcomes. (C-D) Distributions of adjusted (A-B). Adjusted measures are obtained from linear mixed models. The average analysis sample size is shown above each facet. Black dashed lines mark $\ln 3$ and $\ln 3$. The red dashed line marks $\ln 0.05$.

### v3

@@ Abstract
Concerns about the misuse and misinterpretation of $p$-values and statistical significance have motivated alternatives for quantifying evidence. We define a generalized form of Jeffreys’s approximate objective Bayes factor ($\mathit{eJAB}$), a one-line calculation that is a function of the $p$-value, sample size, and dimension of the test. We establish conditions under which $\mathit{eJAB}$ is model-selection consistent and verify them for ten common statistical tests. We assess finite-sample accuracy by comparing $\mathit{eJAB}$ with Markov chain Monte Carlo computed Bayes factors in 12 simulation studies. We then apply $\mathit{eJAB}$ to 71,126 results from ClinicalTrials.gov (CTG) and find that the proportion of findings with $p\text{-value}\leqslant\alpha$ yet $\mathit{eJAB}_{01}>1$ is close to $\alpha$ over a wide range, suggesting that such contradictions are pointing to the type I errors. We catalog 4,088 such candidate type I errors and provide details for 131 with reported $p$-value $\leqslant.01$. We also identify 487 instances of the Jeffreys–Lindley paradox in clinical trials. Finally, we estimate that 75% (6%) of clinical trial plans from CTG set $\alpha\geqslant.05$ as the target evidence threshold, and that 35.5% (0.22%) of results significant at $\alpha=.05$ correspond to evidence that is no stronger than anecdotal under $\mathit{eJAB}$.
@@
Concerns about the replicability of scientific results—stemming in part from the overreliance on null-hypothesis significance testing (NHST) at traditional $p$-value thresholds [1]—have prompted calls for more stringent evidence criteria [2, 3], more judicious interpretation [4], and alternative measures of evidence [5, 6, 7, 8]. The Bayes factor—the ratio of marginal likelihoods under competing hypotheses—can provide evidence for either the null or the alternative [9, 10, 11, 12, 13, 14]. It represents how data update the prior model odds to form the posterior model odds. The evidential behavior of both $p$-values and Bayes factors has been analyzed by stochastic ordering in sample size and effect size, with implications for the Jeffreys–Lindley paradox [15, 16]. Emphasizing effect sizes rather than $p$-value significance thresholds, a quadratic exponential relationship between the Bayes factor and the separation of credible intervals is described in [17]. However, computation remains a practical hurdle: Bayes factors require numerical integration over prior distributions, often via Markov chain Monte Carlo (MCMC) with model-specific tuning [18, 19, 20]. While this computation is feasible with existing algorithms, a simple Bayes factor calculation that requires only the $p$-value, sample size, and dimension of the test has greater applicability to situations where the full dataset is unavailable.
@@
With an emphasis on simplicity and practicality, several approximate Bayes factors based on frequentist calculations have been proposed. For example, certain Bayes factors, with informed priors on a focal parameter, yield fast approximations and are based on Jeffreys’s approximate objective Bayes factor (JAB) [21, 22, 23, 24]. The Bayesian information criterion (BIC) has been extensively used to approximate the Bayes factor [14, 25]. Minimum-Bayes-factor calibrations provide conservative evidence assessments when only a $p$-value is available [26]. Bayes factors based on parametric test statistics are developed by [27], with model-selection consistency under the prior predictive and connections to BIC examined in [28]; extensions to nonparametric statistics appear in [29]. A proposed $3p\sqrt{n}$ rule, based on a piecewise approximation to the $\chi^{2}_{1}$ quantile function, facilitates JAB in the scalar-parameter case [30]. We define $\mathit{eJAB}$, a generalization of JAB that incorporates a finite-sample correction and explicitly accounts for the dimension of the test. It is a one-line calculation that is model-selection consistent under broad conditions. $\mathit{eJAB}$ reduces to JAB when the parameter under test is a scalar and as $n\to\infty$.
@@
Our work differs from [25] in its consideration of $p$-values beyond those from the likelihood ratio test and in our explicit statement of conditions for model-selection consistency. For the latter, [25] demonstrates asymptotic equivalence to the BIC so that model-selection consistency arises under the conditions specified in [31]. It differs from [23, 24, 30] primarily in our formulation of the generalization and in our development of a theoretical justification.
@@
Our contributions are thus threefold. First, we define an extension of JAB, $\mathit{eJAB}$, which generalizes to multidimensional parameters by incorporating the dimension. Our approach differs from BIC-based approximations by accommodating $p$-values beyond those from likelihood-ratio tests, and from other methods by establishing a model-selection consistency justification under regularity conditions. Second, we assess the finite-sample performance of $\mathit{eJAB}$ across ten common parametric and nonparametric tests. The results suggest that $\mathit{eJAB}$ is a broadly applicable and reliable approach to approximate the Bayes factor. Third, we apply $\mathit{eJAB}$ to 71,126 reported findings from https://ClinicalTrials.gov (CTG), accessed on August 7th, 2025, and argue that NHST ($p$-value $\leqslant\alpha$) and $\mathit{eJAB}$ (>1>1, favoring the null) contradictions are candidate type I errors. We flag 4,088 such results and find that the distribution of statistical evidence is uniform across five clinical trial phases for both primary and secondary outcomes in CTG. We also identify 487 instances of the Jeffreys–Lindley paradox.
@@ 1 Methodology
where $q$ is the dimension of the test, $n$ is the sample size, $Q_{\chi^{2}_{q}}(\cdot)$ is the quantile function of the chi-squared distribution with $q$ degrees of freedom (df), and $p$ is the $p$-value from an NHST. We further define $\mathit{eJAB}_{10}=1\ \!/\ \!\mathit{eJAB}_{01}$.
@@
Assuming that the prior density for any nuisance parameters under $\mathcal{H}_{0}$ is everywhere equal to the conditional prior density for those parameters under $\mathcal{H}_{1}$, given the null value of the parameter under test, the Bayes factor (BF) can be expressed as the Savage–Dickey density ratio [32],
@@ 2 Simulation Studies
We used $p$-values from the $t$-test for items 1 and 2; the $F$-test for items 4 and 5; the Wald test for items 3 and 7; and the asymptotic null distribution for item 6. The $p$-values for 8–10 are based on asymptotic distributions. The effective sample size $n$ was taken as the number of uncensored observations for survival analysis and as the number of independent observations (i.e., the product of the number of participants and one less than the number of conditions) for rANOVA [34, 33]. For the other designs, $n$ always refers to the total number of observations or counts. In particular, we tested the harmonic mean of group sizes for item 1, as in JAB [30], but do not demonstrate it here. Among the tests with multiple df ($q>1$) are items 4 through 6, and 10.
Figure 1: Sampling distributions of $D_{n}=-\sqrt{n}\cdot p\ln{p}$ under (A–B) the fixed and (C) the local $\mathcal{H}_{1}$ for eleven statistical tests. Panel labels: ttest ($t$-test), lm (linear regression), glm (logistic regression), anova (analysis of variance), ranova (repeated-measures ANOVA), chisq (chi-squared test), coxph (Cox proportional-hazards regression), wilcox (Wilcoxon signed-rank test), Utest (Mann–Whitney test), Htest (Kruskal–Wallis test), and clogit (conditional logistic regression).
Figure 2: Comparison of the natural logarithmic (A) $\mathit{eJAB}_{10}$ and (B) BIC approximation to the Bayes factor with alternative Bayes-factor computations at a fixed sample size. The $y$-axis shows the MCMC-based Bayes factor for all panels, except for the two chi-squared tests, which use the test-statistic Bayes factor instead. The red text indicates the effective sample size $n$, except for the Cox model, which shows the number of observations. The dashed lines indicate $\mathit{eJAB}_{10}$ of 1/3 and 3, representing moderate evidence against and for an effect, respectively.
@@ 2.1 Verifying the Regularity Conditions
Next, we computed $D_{n}$ of R2 using the $p$-values and sample sizes from the simulated data, generated with medium effect sizes, across the ten designs, plus conditional logistic regression used in our application. In this subsection, items 4 and 5 ($\rho=.9$) each have 41=34-1=3 df; item 6 has $(3-1)\times(3-1)=4$ df under a multinomial design; item 7 represents a semi-parametric Cox model; item 10 has 51=45-1=4 df. Due to randomness, the $x$-axis has to display the total number of observations or pairs, rather than the number of uncensored observations for item 7, or the number of matched pairs for conditional logistic regression, as required in $D_{n}$ and $\mathit{eJAB}$. Condition R2 appears to hold for all fixed alternatives we examined (Fig. 1A–B). For local alternatives of the form $\theta^{*}=\theta_{0}+c/n^{\kappa}$ in the one-sample $t$-test, R2 does not hold at $\kappa=1/2$ but does hold at $\kappa=1/4$ (Fig. 1C). This follows from the denominator of the $t$-statistic being $\mathcal{O}_{p}(1/\sqrt{n})$.
@@ 2.2 Accuracy of the Bayes-Factor Approximation
Pairs of $(\ln\mathit{eJAB}_{10},\ \ln\mathit{BF}_{10})$ are color- and shape-coded by the $p$-value. Overall, $\mathit{eJAB}$ tracks the MCMC-based BFs closely (Fig. 2A). Some inaccuracy, $\ln\mathit{eJAB}_{10}<\ln\mathit{BF}_{10}$, appears for (binary) logistic regression when evidence favors $\mathcal{H}_{1}$. For chi-squared tests, agreement with test-statistic BFs is excellent when evidence favors $\mathcal{H}_{1}$. As evidence shifts toward $\mathcal{H}_{0}$, $\mathit{eJAB}_{10}$ decreases toward zero, as expected under model-selection consistency, whereas the test-statistic BF is inherently bounded below by 1. Additional comparisons with Dirichlet BFs for chi-squared tests are provided in Appendix B. The two Bayes factors reach broad agreement, although differences persist at larger $n$ (Appendix B, Fig. 5). At $n=1,500$, under $\mathcal{H}_{0}$, the Dirichlet $\mathit{BF}_{10}<1/3$ is almost always true, whereas $\mathit{eJAB}_{10}<3$ is mostly observed; under $\mathcal{H}_{1}$, $\mathit{eJAB}_{10}$ more readily identifies an effect. Replacing $\mathit{eJAB}$ with the BIC approximation to the BF performs well when $q=1$, but marked severe discrepancies for $q>1$, indicating that $\mathit{eJAB}$ better approximates BF for ANOVA, rANOVA, and chi-squared tests (Fig. 2B).
@@ 2.3 Bayesian Hypothesis Tests Using Nonparametric Statistics
3 Detecting Type I Errors in 71,126 Clinical Trial Findings with eJAB
Figure 3: Evidence in 71,126 results. (A–B) Proportion of analyses with $p\leqslant\alpha$ and $\mathit{eJAB}_{01}>1$ as candidates type I errors, with $p=.05$ excluded in (B). The red dashed line marks the overall share with $\mathit{eJAB}_{01}>1$ as $\alpha\rightarrow 1$. (C) Proportion of $p\leqslant.05$ with Bayes-factor threshold $K$. The inset highlights the Jeffreys–Lindley paradox region. (D) Scatterplot of $\ln\mathit{eJAB}_{01}$ versus $p$-value, with an inset zoom for $p\leqslant.05$. Shaded bands denote evidence categories. (E) Posterior distribution of the proportion of $p\leqslant.05$ and $\mathit{eJAB}_{01}>1/3$. The red dashed line indicates the posterior mean. (F) Density of sample sizes for all $p\leqslant.05$, stratified by Bayes-factor category.
@@ 3 Detecting Type I Errors in 71,126 Clinical Trial Findings with eJAB
Our objective is to characterize the overall strength of evidence in 71,126 CTG findings and to identify potential type I errors. These results span 43 MeSH categories (e.g., neoplasms, cardiovascular diseases, and autoimmune diseases) and encompass nine statistical tests: one-sample $t$-test, two-sample $t$-test, linear regression, logistic regression, ANOVA, conditional logistic regression, Wilcoxon signed-rank test, Mann–Whitney test, and Kruskal–Wallis test. We verify that regularity condition R2 holds for conditional logistic regression (Fig. 1C).
@@
We therefore classify findings with $p\leqslant\alpha$ and $\mathit{eJAB}_{01}>1$ as candidate type I errors at the significance level $\alpha$. Among 30,790 results significant at $\alpha=.05$, we catalog 4,088 candidates. Although lower thresholds have been recommended in the literature [2, 3], $\alpha=.05$ remains germane in CTG: in a random sample of 52 trial protocols, 75% (SE 6%) explicitly set $\alpha\geqslant.05$ as the target evidential threshold. We list all 4,088 candidates on the Open Science Framework at https://osf.io/qeprj/ (see ‘Supplementary Information’) and provide an annotated subset of 131 cases with $p\leqslant.01$ and $\mathit{eJAB}_{01}>1$, cross-referenced with CTG and PubMed.
@@
We also examine the proportion of all $p\leqslant.05$ and $\mathit{eJAB}_{01}\geqslant K$ for $K\in(0,300)$; at $K=1$, this proportion recovers the candidate type I error at $\alpha=.05$. The inset highlights Jeffreys–Lindley paradox cases (Fig. 3C). We identify 487 such instances in the CTG data. Additional results stratified by MeSH category appear in Appendix B (Figs. 68).
@@
We examine the distributions of $\ln\mathit{eJAB}_{01}$ and $\ln{p}$ by trial phase for primary and secondary outcomes (Fig. 4A–B). For primary outcomes, there is no discernible trend toward stronger evidence from early phase to phase 4. The distributions of $\ln{p}$ shift downward at phase 3, consistent with the markedly larger average sample sizes at that phase. For secondary outcomes, we again observe no phase-wise increase in evidence, though variability in both measures is higher in phase 3. Plotting $p$ and sample size by phase shows that changes in $p$ track changes in sample size (Appendix B, Fig. 10).
To better visualize trends in evidence across phases, we fit a linear mixed model with a fixed intercept and trial-specific random effects to each of $\ln\mathit{eJAB}_{01}$ and $\ln{p}$, and then examine residuals re-centered by the estimated intercept (Appendix B, Fig. 9). The adjusted measures are essentially uniform across phases (Fig. 4C–D). For primary outcomes, most adjusted $p$-values lie near .05 and above .01. Using $\mathit{eJAB}$, this interpretation corresponds to ‘anecdotal evidence’ or ‘evidence that is barely worth mentioning’ for both primary and secondary outcomes [42, p. 105]. Comparing primary with secondary outcomes, we note a slight and phase-invariant shift toward greater evidence for the alternative in primary outcomes.
Figure 4: Measures of evidence by clinical trial phase. (A–B) Distributions of $\ln\mathit{eJAB}_{01}$ and $\ln{p}$ by phase for primary (A) and secondary (B) outcomes. (C–D) Distributions of adjusted (A–B). Adjusted measures are obtained from linear mixed models. The average analysis sample size is shown above each facet. Black dashed lines mark $\ln 3$ and $\ln 3$. The red dashed line marks $\ln 0.05$.